# Data-adaptive structural change-point detection via isolation

### v2

@@ 5 Simulations
Proceeding now with the piecewise-linear framework, the simulation results in Tables 5 and 6 show that in the case of piecewise-linear signals, DAIS performs at least as accurately as the algorithms ID_ic and CPOP, and better than other competitors available in the literature regarding the estimated number of change-points and their locations. The average computational time for DAIS is very low, and most of the times, significantly lower than the best-performing competitors. For signal (S10), DAIS outperforms all competitors except for CPOP in terms of the accuracy for the number of change-points detected. However, CPOP has a higher MSE and is at least four times slower compared to DAIS in all the examples presented.
@@
Table 3: Distribution of $\hat{N}-N$ over 100 simulated data sequences of the signals (S1), (S2), (S3) and (S4). The average MSE, $d_{H}$ and computational times are also given.
@@
DAIS 0 15 0 81 2 2 0 0.014 0.089 0.004
ID_th 0 8 0 72 9 10 1 0.014 0.098 0.004
ID_ic 0 24 0 73 3 0 0 0.016 0.133 0.002
WBS_th 0 0 5 30 18 25 22 0.018 0.196 0.009
WBS_ic 0 30 0 68 2 0 0 0.016 0.166 0.025
WBS2 (S1) 0 10 2 75 7 1 5 0.015 0.091 0.446
PELT 0 77 0 23 0 0 0 0.026 0.397 0.001
NOT 0 26 0 72 2 0 0 0.016 0.146 0.024
MOSUM 0 15 0 74 6 4 1 0.014 0.108 0.006
MSCP 0 75 23 2 0 0 0 0.030 0.394 2.600
SeedBS_th 0 0 4 20 21 22 33 0.019 0.245 0.013
SeedBS_ic 0 32 0 66 2 0 0 0.017 0.176 0.013
DP_univar 0 0 0 6 3 10 81 0.045 0.334 0.138
@@
DAIS 0 0 0 98 2 0 0 0.253 0.004 0.0014
ID_th 0 0 0 96 4 0 0 0.254 0.006 0.0008
ID_ic 22 11 42 24 1 0 0 1.540 0.111 0.0013
WBS_th 0 0 0 87 8 4 1 0.251 0.007 0.0009
WBS_ic 0 0 0 71 19 9 1 0.270 0.012 0.0018
WBS2 (S4) 0 0 0 99 1 0 0 0.235 0.004 0.0018
PELT 4 4 0 92 0 0 0 0.380 0.018 0.0001
NOT 0 0 0 92 7 1 0 0.242 0.007 0.0144
MOSUM 100 0 0 0 0 0 0 6.460 0.893 0.0001
MSCP 100 0 0 0 0 0 0 6.270 0.446 0.0257
SeedBS_th 0 0 3 86 6 5 0 0.267 0.008 0.0108
SeedBS_ic 0 0 0 73 20 5 2 0.265 0.011 0.0108
DP_univar 100 0 0 0 0 0 0 5.790 0.154 0.0001
@@
MSCP 100 0 0 6.220 0.160 1.090
SeedBS_th 0 0 100 0.235 0.003 0.012
SeedBS_ic 100 0 0 6.250 0.992 0.013
DP_univar 100 0 0 5.710 0.025 0.030
@@
Method Signal $-1$ 0 1 $2$ $\geq 3$ MSE $d_{H}$ Time (s)
DAIS 0 99 1 0 0 0.030 0.082 0.024
ID_th 0 71 28 1 0 0.036 0.130 0.024
ID_ic 0 95 5 0 0 0.031 0.096 0.008
CPOP 0 99 1 0 0 0.013 0.050 5.550
NOT (S8) 0 99 1 0 0 0.016 0.060 0.186
MARS 0 5 39 47 9 0.025 0.194 0.008
TF 0 0 0 0 100 0.018 0.440 0.768
TS 1 99 0 0 0 0.096 0.186 1.020
local_poly 0 1 5 12 82 0.054 0.503 125.400
@@ 6.2 Heavy-tailed noise
$X^{\ast}_{q}=\frac{1}{s}\sum_{t=(q-1)s+1}^{qs}X_{t},\text{ for }t=1,2,\ldots,q\quad\text{and}\quad X^{\ast}_{Q}=\frac{1}{T-(Q-1)s}\sum_{t=(Q-1)s+1}^{T}X_{t}.$
@@
The performance of DAIS in the case of heavy-tailed data can be found in Table 7. For the pre-averaging technique we use $s=5$ and the heavy-tailed noise is chosen to be the Student’s t-distribution with degrees of freedom being equal to either 5 or 7. Signals (S1) and (S7) are piecewise-constant and (S8) is piecewise-linear and their description can be found in Appendix A. As expected, the accuracy of DAIS is reduced compared to Section 5. Moreover, the more heavy-tailed the noise, the more the overdetection. This is again expected, as the spikes caused by the heavy-tailed distribution of $\epsilon_{t}$ can be mistakenly detected as change-points by the algorithm, even after the pre-processing. However, DAIS performs relatively well for signal (S1) even for 5 degrees of freedom.
@@ 6.3 Multivariate models
where $\boldsymbol{X_{t}}\in\mathbb{R}^{d\times 1}$ are the observed data and $\boldsymbol{f_{t}}\in\mathbb{R}^{d\times 1}$ is the d-dimensional deterministic signal with structural changes at certain points. In the case of piecewise-constant signals, the structure of $\boldsymbol{f_{t}}$ is given by $\boldsymbol{f_{t}}=\boldsymbol{\mu_{j}}$ for $t\in\{r_{j-1}+1,\ldots,r_{j}\}$ and $\boldsymbol{f_{r_{j}}}\neq\boldsymbol{f_{r_{j}+1}}$ where $\boldsymbol{\mu_{j}}\in\mathbb{R}^{d\times 1}$ for $j=1,2,\ldots,N+1$. For continuous piecewise-linear signals, $\boldsymbol{f_{t}}=\boldsymbol{\mu_{j,1}}+\boldsymbol{\mu_{j,2}}t$ for $t\in\{r_{j-1}+1,\ldots,r_{j}\}$ and $\boldsymbol{f_{r_{j}-1}}+\boldsymbol{f_{r_{j}+1}}\neq 2\boldsymbol{f_{r_{j}}}$ where $\boldsymbol{\mu_{j,1}},\boldsymbol{\mu_{j,2}}\in\mathbb{R}^{d\times 1}$. To ensure continuity, we need the additional constraint of $\boldsymbol{\mu_{k,1}}+\boldsymbol{\mu_{k,2}}r_{k}=\boldsymbol{\mu_{k+1,1}}+\boldsymbol{\mu_{k+1,2}}r_{k}$ for $k\in\{1,2,\ldots,N\}$. As with the univariate case, the algorithm can also be applied to more general signal structures.
@@
We define $\boldsymbol{y_{t}}$ to be the $\mathbb{R}^{d\times 1}$ vector with entries the value of the chosen contrast function, this being (10) for piecewise-constant signals and (12) for piecewise-linear ones. At each step of the algorithm, the goal is to decide whether the point that maximizes $L(\boldsymbol{y_{t}})$ is a change-point. The aggregation scheme for the contrast function, as just described, was also proposed by [3]. Algorithm 2 provides a pseudocode for DAIS in the case of multivariate signals, which we call MDAIS. For a discussion on the choice of the mean-dominant norm, we refer the reader to [1, 3, 12, 15].
Theoretical results for this variant of the algorithm about consistency for the number and locations of the estimated change-points can be proven. In the rest of this section, we provide the consistency theorem that concerns the case of a piecewise-constant $\boldsymbol{f_{t}}$ and the outline of its proof can be found in Appendix E. However, since the focus of this paper is proposing the novel data-adaptive methodology, we do not delve further in this. Following a similar methodology, theoretical results can also be proved for $\boldsymbol{f_{t}}$ being continuous piecewise-linear. For the statement of the theorem, we require the following assumption, which is equivalent to assumption (A2) for univariate $f_{t}$:
@@
Similarly to the univariate case, the number of change-points $N$ is allowed to grow with $T$ and $d$. Note that the threshold constant, $\zeta_{T,d}$ depends on both the length of the data sequence and its dimension.
@@
Let $\{\boldsymbol{X_{t}}\}_{t=1,\ldots,T}$ follow model (20) with $\boldsymbol{f_{t}}$ piecewise-constant and let $\boldsymbol{\epsilon_{t}}\sim\mathcal{N}_{d}(\boldsymbol{0},\Sigma)$, where $\Sigma\in\mathbb{R}^{d\times d}$ is positive definite and also that assumptions (A1) and (A4) hold. Let $N$ and $r_{j},j=1,2,\ldots N$ be the number and location of the change-points, while $\hat{N}$ and $\hat{r}_{j},j=1,2,\ldots\hat{N}$ their estimates, sorted in increasing order. In addition, $\boldsymbol{\Delta_{j}}:=\lvert\boldsymbol{f_{r_{j}+1}}-\boldsymbol{f_{r_{j}}}\rvert,j=1,2,\ldots N$ is the magnitude of each discontinuity in $\boldsymbol{f_{t}}$, $\underline{f}_{T}:=\inf_{j=1,\ldots,N}\{L\left(\boldsymbol{\Delta_{j}}\right)\}$ and $\delta_{T}=\min_{j=1,2,\ldots,N+1}\lvert r_{j}-r_{j-1}\rvert$, where $r_{0}=0,r_{N+1}=T$. Then there exist positive constants $C_{1},C_{2},C_{3},C_{4}$ which do not depend on $T$ such that for $C_{1}\sqrt{\log\left(Td^{1/4}\right)}\leq\zeta_{T,d}<C_{2}\sqrt{\delta_{T}}\underline{f}_{T}$ and for sufficiently large $T$, we obtain
$\mathbb{P}\Biggl{(}\hat{N}=N,\max_{j=1,2,\ldots,N}\biggl{\{}\left|\hat{r}_{j}-r_{j}\right|\alpha^{2}\biggl{\}}\leq C_{3}\log\left(Td^{1/4}\right)\Biggl{)}\geq 1-\frac{C_{4}}{T}.$ (23)
where $\alpha=\Delta^{q_{j}}_{j}=\lvert f_{r_{j+1},q_{j}}-f_{r_{j},q_{j}}\rvert$ when $L(\cdot)=L_{\infty}(\cdot)$, and $\alpha=L_{2}\left(\boldsymbol{\Delta_{j}}\right)$ when $L(\cdot)=L_{2}(\cdot)$, $q_{j}:=\textrm{argmax}_{k=1,\ldots,d}\tilde{X}^{\hat{r}_{j},k}_{s_{j},e_{j}}$ for $[s_{j},e_{j}]$ being the interval where $\hat{r}_{j}$ is obtained and $\tilde{X}_{s,e}^{b,j}=\sqrt{\frac{e-b}{\ell(b-s+1)}}\sum_{t=s}^{b}X_{t,j}-\sqrt{\frac{b-s+1}{\ell(e-b)}}\sum_{t=b+1}^{e}X_{t,j}$, $\ell=e-s+1$.
The outline of the proof of Theorem 3 is given in Appendix E. Note that this extension of our algorithm is robust to spatial dependence. Theoretical results related to this statement, as well as simulation results can be found in [3], where the extension of ID ([2]), namely the MID algorithm, is introduced.
@@ 7.1 Crime data
In this section, DAIS is applied to real data and the underlying signal is assumed to be piecewise-constant. The chosen dataset includes daily crime reports in Montgomery County, Maryland and can be found in https://catalog.data.gov/dataset/crime. We use daily observations starting from 22/01/2020, up to 31/08/2024. Since the data involve counts of crimes, they are positive integer-valued and thus a transformation is required to bring them closer to Gaussian data with constant variance. This is done using the Anscombe transform, [4], $\alpha:\mathbb{N}\rightarrow\mathbb{R}$, with $\alpha(x)=2\sqrt{x+3/8}$.
@@
The top plot of Figure 3 is a plot of the real data (in black), along with the estimated underlying signal, $\hat{f}_{t}$, according to the change-points detected by DAIS (plotted in red). We now attempt to provide a possible explanation about the change-points estimated by DAIS that express the most important movements in the data. The first change-point occurs on the $17^{\textrm{th}}$ of March 2020 and corresponds to the first days of positive cases of COVID-19 when the first rules were imposed to limit the spread of the virus among the community in the state of Maryland. As expected, since people were forced to stay at home, the number of crimes reported dropped significantly. The second change-point is on the $28^{\textrm{th}}$ of April 2020, when the official authorities started lifting the restrictions and the number of crimes increased. Some small decrease around the $11^{\textrm{th}}$ of December 2020 can be explained by people’s behavioural change around the holidays. The number of crimes returns close to the previous levels, on the $28^{\textrm{th}}$ of May 2021 and a further increase is observed at the end of the summer, $27^{\textrm{th}}$ of August 2021. This could possibly be due to people going on holidays out of town, leaving their homes more vulnerable to break-ins. As has also been observed for 2020, there is again a reduced number of reported crimes around the Christmas holidays of 2021, starting from the $23^{\textrm{th}}$ of December 2021 until the $17^{\textrm{th}}$ of January 2022. It is notable that the next change-point detected is almost a year later, as there is a sudden decrease on the $23^{\textrm{rd}}$ of December 2022, with an increase on the $26^{\textrm{th}}$ of the same month. An increase in the daily number of crimes reported is also detected on the $28^{\text{th}}$ of September 2023. As in the previous years, there is a drop on the $23^{\textrm{rd}}$ of December 2023 and an increase to a slightly higher average than before on the $23^{\textrm{rd}}$ of January 2024. Further increase to the number of daily crimes reported is observed on the $26^{\textrm{th}}$ of February and the $10^{\textrm{th}}$ of April 2024, with a decrease on the $4^{\textrm{th}}$ of June 2024. It is clear that the average number of crimes has increased in the last years. It is important to note that there is a clear trend as change-points were detected by DAIS on 11/12/2020, 23/12/2021, 23/12/2022 and 23/12/2023. All four are near the Christmas holidays and a decrease in the daily number of crimes reported is noticed, followed by the trend increasing on the 17/01/2022, 26/12/2022 and 22/01/2024.
@@
The fits obtained by three competitors, ID, MOSUM, and NOT, are shown in the bottom plot of Figure 3, plotted with blue, yellow and green lines, respectively. All methods obtain different signals on the same data, with ID and DAIS providing similar fits and NOT having minor differences with them. ID and NOT detect 11 and 10 out of the 15 change-points that DAIS detects, respectively. They both don’t detect the change-points in May 2021 and the last 2 change-points detected by DAIS, in April and June of 2024. ID also does not detect the change-point just before Christmas 2020, while NOT does not detect the 2 change-points around Christmas 2022. MOSUM detects only 5 change-points, with the only change-point around the Christmas period being the one in 2020, which ID missed. By detecting change-points around the same period every year, we could argue that DAIS is detecting the seasonality around the Christmas holidays, which none of the other competitors managed to fully capture.
@@
Figure 3: The observed time series data of daily crime reports in Montgomery County and fitted piecewise-constant mean signal obtained by applying DAIS (red), ID (blue), MOSUM (yellow) and NOT (green) to the data. The x-axis labels correspond to the month and year.

### v3

@@ 5 Simulations
We now proceed with a discussion of the performance of DAIS under the continuous, piecewise-linear framework. The simulation results are shown in Tables 6 and 7. We conclude that DAIS performs at least as accurately as the algorithms ID and CPOP, and substantially better than other competitors available in the literature, regarding accuracy in terms of both the estimated number and the estimated locations of the change-points. Furthermore, the average computational time for DAIS remains very low in all signals tested; in fact, it is the lowest among the times for the top performing methods.
$\hat{N}-N$
Method Signal $-2$ $-1$ 0 1 2 $\geq 3$ MSE $d_{H}$ Time (s)
DAIS 16 0 80 2 2 0 0.014 0.089 0.004
ID_th 10 0 79 8 3 0 0.013 0.081 0.007
ID_ic 33 0 64 2 1 0 0.018 0.179 0.013
WBS_th 8 7 63 16 5 1 0.015 0.105 0.010
WBS_ic 30 0 68 2 0 0 0.016 0.166 0.032
WBS2 10 2 75 7 1 5 0.015 0.091 0.563
PELT 77 0 23 0 0 0 0.026 0.397 0.001
NOT (S1) 30 0 68 2 0 0 0.016 0.167 0.037
MOSUM 15 0 74 6 4 1 0.014 0.108 0.011
MSCP 75 23 2 0 0 0 0.030 0.394 2.600
SeedBS_th 11 18 55 13 2 1 0.017 0.103 0.022
SeedBS_ic 32 0 66 2 0 0 0.017 0.176 0.022
DP_univar 0 0 6 3 10 81 0.045 0.334 0.138
SMUCE 4 36 53 7 0 0 0.020 0.057 0.005
MsFPOP 29 0 70 1 0 0 0.016 0.157 0.004
DAIS 0 7 86 4 3 0 0.063 0.039 0.001
ID_th 0 37 49 10 2 2 0.125 0.065 0.001
ID_ic 0 9 83 6 1 1 0.068 0.026 0.004
WBS_th 0 1 64 17 8 10 0.075 0.154 0.002
WBS_ic 0 4 86 7 2 1 0.061 0.025 0.005
WBS2 0 24 67 3 3 3 0.087 0.048 0.055
PELT 0 18 81 1 0 0 0.068 0.011 0.001
NOT (S2) 0 4 87 6 2 1 0.062 0.024 0.013
MOSUM 0 83 14 3 0 0 0.185 0.050 0.002
MSCP 0 100 0 0 0 0 0.197 0.069 0.066
SeedBS_th 0 5 75 12 7 1 0.064 0.073 0.026
SeedBS_ic 0 6 87 6 1 0 0.060 0.021 0.026
DP_univar 0 15 59 13 11 2 0.183 0.175 0.001
SMUCE 0 36 59 5 0 0 0.111 0.033 0.001
MsFPOP 0 12 86 2 0 0 0.066 0.015 0.001
@@
Table 4: Distribution of $\hat{N}-N$ over 100 simulated data sequences of the Signals (S3), (S4) and (S5). The average MSE, $d_{H}$ and computational times are also given.
@@
SMUCE 0 0 2 98 0 0 0 1.560 0.008 0.002
MsFPOP 0 0 0 0 0 0 100 11.80 0.153 0.001
DAIS 0 0 0 98 2 0 0 0.253 0.004 0.001
ID_th 0 0 0 99 1 0 0 0.246 0.005 0.003
ID_ic 0 0 0 91 9 0 0 0.259 0.007 0.004
WBS_th 0 0 0 94 6 0 0 0.239 0.005 0.001
WBS_ic 0 0 0 71 19 9 1 0.270 0.012 0.003
WBS2 0 0 0 99 1 0 0 0.235 0.004 0.024
PELT 4 4 0 92 0 0 0 0.380 0.018 0.001
NOT (S5) 0 0 0 72 17 8 3 0.271 0.012 0.004
MOSUM 100 0 0 0 0 0 0 6.460 0.893 0.001
MSCP 100 0 0 0 0 0 0 6.270 0.446 0.026
SeedBS_th 0 1 14 80 5 0 0 0.348 0.016 0.019
SeedBS_ic 0 0 0 73 20 5 2 0.265 0.011 0.019
DP_univar 100 0 0 0 0 0 0 5.790 0.154 0.001
SMUCE 14 10 19 57 0 0 0 1.110 0.034 0.001
MsFPOP 0 0 0 100 0 0 0 0.228 0.004 0.001
@@
MSCP 99 1 0 6.220 0.160 1.090
SeedBS_th 0 0 100 0.402 0.007 0.023
SeedBS_ic 100 0 0 6.250 0.992 0.023
DP_univar 0 100 0 5.710 0.025 0.030
SMUCE 0 100 0 5.430 0.020 0.002
MsFPOP 0 0 100 0.209 0.002 0.001
@@
Method Signal $\leq-2$ $-1$ 0 1 $2$ $\geq 3$ MSE $d_{H}$ Time (s)
DAIS 0 0 99 1 0 0 0.030 0.085 0.020
ID_th 0 0 95 5 0 0 0.030 0.086 0.024
ID_ic 0 0 93 7 0 0 0.034 0.099 0.093
CPOP 0 0 99 1 0 0 0.013 0.050 5.550
NOT (S9) 0 0 100 0 0 0 0.015 0.055 0.720
MARS 0 0 5 39 47 9 0.025 0.194 0.008
TF 0 0 0 0 0 100 0.018 0.440 0.768
TS 0 1 99 0 0 0 0.096 0.186 1.020
local_poly 0 0 1 5 12 82 0.054 0.503 125.400
@@ 6.2 Heavy-tailed noise
$X^{\ast}_{q}=\frac{1}{s}\sum_{t=(q-1)s+1}^{qs}X_{t},\text{ for }{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}q=1,2,\ldots,Q-1}\quad\text{and}\quad X^{\ast}_{Q}=\frac{1}{T-(Q-1)s}\sum_{t=(Q-1)s+1}^{T}X_{t}.$
@@
The performance of DAIS in the case of heavy-tailed data can be found in Table 8. For the pre-averaging technique we use $s=5$ and the heavy-tailed noise is chosen to be the Student’s t-distribution with degrees of freedom being equal to either 5 or 7. Signals (S1) and (S8) are piecewise-constant and (S9) is piecewise-linear and their description can be found in Appendix A. Even though the accuracy of DAIS is reduced compared to Section 5, the results indicate a good level of robustness for DAIS at the presence of heavy-tailed noise. The more heavy-tailed the noise, the more the overdetection, which is expected, as spikes caused by the heavy-tailed distribution of $\epsilon_{t}$ can sometimes be mistakenly detected as change-points. However, DAIS performs relatively well for Signals (S1) and (S8) even for the case of 5 degrees of freedom for the Student’s t-distribution.
@@ 6.3 Multivariate models
where $\boldsymbol{X_{t}}\in\mathbb{R}^{d\times 1}$ are the observed data and $\boldsymbol{f_{t}}\in\mathbb{R}^{d\times 1}$ is the d-dimensional deterministic signal with structural changes at certain points. In the case of piecewise-constant signals, the structure of $\boldsymbol{f_{t}}$ is given by $\boldsymbol{f_{t}}=\boldsymbol{\mu_{j}}$ for $t\in\{r_{j-1}+1,\ldots,r_{j}\}$ and $\boldsymbol{f_{r_{j}}}\neq\boldsymbol{f_{r_{j}+1}}$ where $\boldsymbol{\mu_{j}}\in\mathbb{R}^{d\times 1}$ for $j=1,2,\ldots,N+1$. For continuous, piecewise-linear signals, $\boldsymbol{f_{t}}=\boldsymbol{\mu_{j,1}}+\boldsymbol{\mu_{j,2}}t$ for $t\in\{r_{j-1}+1,\ldots,r_{j}\}$ and $\boldsymbol{f_{r_{j}-1}}+\boldsymbol{f_{r_{j}+1}}\neq 2\boldsymbol{f_{r_{j}}}$ where $\boldsymbol{\mu_{j,1}},\boldsymbol{\mu_{j,2}}\in\mathbb{R}^{d\times 1}$. To ensure continuity, we need the additional constraint of $\boldsymbol{\mu_{k,1}}+\boldsymbol{\mu_{k,2}}r_{k}=\boldsymbol{\mu_{k+1,1}}+\boldsymbol{\mu_{k+1,2}}r_{k}$ for $k\in\{1,2,\ldots,N\}$. As with the univariate case, the algorithm can also be applied to more general signal structures.
@@
We define $\boldsymbol{y_{t}}$ to be the $\mathbb{R}^{d\times 1}$ vector with entries the value of the chosen contrast function, this being (15) for piecewise-constant signals and (17) for piecewise-linear ones. At each step of the algorithm, the goal is to decide whether the point that maximizes the mean-dominant norm $L(\boldsymbol{y_{t}})$ is a change-point. The aggregation scheme for the contrast function, as just described, was also proposed by [4]. Algorithm 2 provides a pseudocode for DAIS in the case of multivariate signals, which we call MDAIS. For a discussion on the choice of the mean-dominant norm, we refer the reader to [2, 4, 14, 17].
Theoretical results for this variant of the algorithm about consistency for the number and locations of the estimated change-points can be proven. In the rest of this section, we provide the consistency theorem that concerns the case of a piecewise-constant $\boldsymbol{f_{t}}$ and the outline of its proof can be found in Appendix C. However, since the focus of this paper is proposing the novel data-adaptive methodology, we do not delve further in this. Following a similar methodology, theoretical results can also be proven for $\boldsymbol{f_{t}}$ being continuous, piecewise-linear. For the statement of the theorem, we require the following assumption, which is equivalent to Assumption (A2) (used for the case of univariate $f_{t}$):
@@
Similarly to the univariate case, the number of change-points $N$ is allowed to grow with $T$ and $d$. Note that the threshold constant, $\zeta_{T,d}$ depends on both the length of the data sequence and its dimension. Before providing the statement of the theorem, we generalize the contrast function chosen for the case of piecewise-constant signals, which is the absolute value of the CUSUM statistic as given in (15), to
$\left|\tilde{X}_{s,e}^{b,j}\right|=\left|\sqrt{\frac{e-b}{\ell(b-s+1)}}\sum_{t=s}^{b}X_{t,j}-\sqrt{\frac{b-s+1}{\ell(e-b)}}\sum_{t=b+1}^{e}X_{t,j}\right|$ (28)
for $\ell=e-s+1$, $b\in\{s,\ldots,e-1\}$ and $j\in\{1,\ldots,d\}$.
@@
Let $\{\boldsymbol{X_{t}}\}_{t=1,\ldots,T}$ follow model (25) with $\boldsymbol{f_{t}}$ piecewise-constant and let $\boldsymbol{\epsilon_{t}}\sim\mathcal{N}_{d}(\boldsymbol{0},\Sigma)$, where $\Sigma\in\mathbb{R}^{d\times d}$ is positive definite and also that Assumptions (A1) and (A4) hold. Let $N$ and $r_{j},j=1,2,\ldots N$ be the number and location of the change-points, while $\hat{N}$ and $\hat{r}_{j},j=1,2,\ldots\hat{N}$ their estimates, sorted in increasing order. In addition, $\boldsymbol{\Delta_{j}}:=\lvert\boldsymbol{f_{r_{j}+1}}-\boldsymbol{f_{r_{j}}}\rvert$ for $j=1,2,\ldots N$, $\underline{f}_{T}:=\inf_{j=1,\ldots,N}\{L\left(\boldsymbol{\Delta_{j}}\right)\}$ and $\delta_{T}$ is as in (14). Then there exist positive constants $C_{1},C_{2},C_{3},C_{4}$ independent of $T$ and $d$, such that for $C_{1}\sqrt{\log\left(Td^{1/4}\right)}\leq\zeta_{T,d}<C_{2}\sqrt{\delta_{T}}\underline{f}_{T}$ and for sufficiently large $T$, we obtain
$\mathbb{P}\Biggl{(}\hat{N}=N,\max_{j=1,2,\ldots,N}\biggl{\{}\left|\hat{r}_{j}-r_{j}\right|\alpha_{j}^{2}\biggl{\}}\leq C_{3}\log\left(Td^{1/4}\right)\Biggl{)}\geq 1-\frac{C_{4}}{T},$ (29)
where, for $\left|\tilde{X}_{s,e}^{b,j}\right|$ as in (28) and $[s_{j},e_{j}]$ the interval where $\hat{r}_{j}$ is obtained, we denote by $q_{j}:={\rm{argmax}}_{k=1,\ldots,d}\left|\tilde{X}^{\hat{r}_{j},k}_{s_{j},e_{j}}\right|$ and
$\alpha_{j}=\left\{\begin{array}[]{cc}\lvert f_{r_{j+1},q_{j}}-f_{r_{j},q_{j}}\rvert,&\textrm{for }L(\cdot)=L_{\infty}(\cdot)\\ L^{2}\left(\boldsymbol{\Delta}_{j}\right),&\textrm{for }L(\cdot)=L_{2}(\cdot)\end{array}\right..$
The outline of the proof of Theorem 3 is given in Appendix C. Note that this extension of our algorithm is robust to spatial dependence. Theoretical results related to this statement, as well as simulation results can be found in [4], where the extension of ID ([3]), namely the MID algorithm, is introduced.
@@ 7.1 Crime data
In this section, DAIS is applied to real data and the underlying signal is assumed to be piecewise-constant. The chosen dataset includes daily crime reports in Montgomery County, Maryland and can be found in https://catalog.data.gov/dataset/crime. We use daily observations starting from 22/01/2020, up to 31/08/2024. Since the data involve counts of crimes, they are positive integer-valued and thus a transformation is required to bring them closer to Gaussian data with constant variance. This is done using the Anscombe transform ([5]) $\alpha:\mathbb{N}\rightarrow\mathbb{R}$, with $\alpha(x)=2\sqrt{x+3/8}$.
@@
The top plot of Figure 3 is a plot of the real data (in black), along with the estimated underlying signal, $\hat{f}_{t}$, according to the change-points detected by DAIS (plotted in red). We now attempt to provide a possible explanation about the change-points estimated by DAIS that express the most important movements in the data. The first change-point occurs on the $17^{\textrm{th}}$ of March 2020 and corresponds to the first days of positive cases of COVID-19 when the first rules were imposed to limit the spread of the virus among the community in the state of Maryland. As expected, since people were forced to stay at home, the number of crimes reported dropped significantly. The second change-point is on the $28^{\textrm{th}}$ of April 2020, when the official authorities started lifting the restrictions and the number of crimes increased. Some small decrease around the $11^{\textrm{th}}$ of December 2020 can be explained by people’s behavioral change around the holidays. The number of crimes returns close to the previous levels, on the $28^{\textrm{th}}$ of May 2021 and a further increase is observed at the end of the summer, $27^{\textrm{th}}$ of August 2021. This could possibly be due to people going on holidays out of town, leaving their homes more vulnerable to break-ins. As has also been observed for 2020, there is again a reduced number of reported crimes around the Christmas holidays of 2021, starting from the $23^{\textrm{th}}$ of December 2021 until the $17^{\textrm{th}}$ of January 2022. It is notable that the next change-point detected is almost a year later, as there is a sudden decrease on the $23^{\textrm{rd}}$ of December 2022, with an increase on the $26^{\textrm{th}}$ of the same month. An increase in the daily number of crimes reported is also detected on the $28^{\text{th}}$ of September 2023. As in the previous years, there is a drop on the $23^{\textrm{rd}}$ of December 2023 and an increase to a slightly higher average than before on the $23^{\textrm{rd}}$ of January 2024. Further increase to the number of daily crimes reported is observed on the $26^{\textrm{th}}$ of February and the $10^{\textrm{th}}$ of April 2024, with a decrease on the $4^{\textrm{th}}$ of June 2024. It is clear that the average number of crimes has increased in the last years. It is important to note that there is a clear trend as change-points were detected by DAIS on 11/12/2020, 23/12/2021, 23/12/2022 and 23/12/2023. All four are near the Christmas holidays and a decrease in the daily number of crimes reported is noticed, followed by the trend increasing on the 17/01/2022, 26/12/2022 and 22/01/2024.
@@
The fits obtained by three competitors, ID, MOSUM, and NOT, are shown in the bottom plot of Figure 3, plotted with blue, yellow and green lines, respectively. All methods obtain different signals on the same data, with ID and DAIS providing similar fits and NOT having minor differences with them. ID and NOT detect 11 and 10 out of the 15 change-points that DAIS detects, respectively. They both do not detect the change-points in May 2021 and the last 2 change-points detected by DAIS, in April and June of 2024. ID also does not detect the change-point just before Christmas 2020, while NOT does not detect the 2 change-points around Christmas 2022. MOSUM detects only 5 change-points, with the only change-point around the Christmas period being the one in 2020, which ID missed. By detecting change-points around the same period every year, we could argue that DAIS is detecting the seasonality around the Christmas holidays, which none of the other competitors managed to fully capture.
@@
Figure 3: The observed time series data of daily crime reports in Montgomery County and fitted piecewise-constant mean signal obtained by applying DAIS (red), ID (blue), MOSUM (yellow) and NOT (green) to the data. The x-axis labels correspond to the month and year in an MM/YY format.